For a convex 51-gon, what is the sum of the measures of the interior angles?
Answer by stanbon(75887)

For a convex 51-gon, what is the sum of the measures of the interior angles?
-----
sum of interior angles = (51-2)180 = 49*180 = 8820 degrees
===================
Cheers,
Stan H.